# 1st Dimitrios Bouzoulas    2nd Eerik Alamikkotervo    3rd Risto Ojala

### v2

@@ I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian-related traffic accidents [1]. According to the World Health Organization (WHO) [2], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors, such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions, nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [3] on the data provided by the transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
Figure 1: Autolabelling pipeline, using unlabelled low-light IR-RGB image pairs as an input, the system outputs labelled RGB images.
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in low-light conditions, as RGB model accuracy is highly affected by lighting [4]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared (IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal, we leverage the lighting invariance of infrared cameras as visible light has minimal effect on infrared radiation.
In our approach, we utilise the benefits of infrared images to detect pedestrians in infrared and then transfer the labels to the corresponding RGB image. Our proposed approach relies on aligned infrared and RGB images, which can be conveniently collected with a research vehicle equipped with a suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work.
The proposed pipeline has not previously been implemented for autolabelling tasks and shows a range of positive initial results. The autolabels generated using the proposed pipeline performed on par or outperformed the ground truth labels in many cases in key metrics such as map@50, map@50-95, and LAMR.
@@ II-A Pedestrian Detection
Early pedestrian detection approaches utilize hand-crafted features such as Histogram of Oriented Gradients (HOG) [5]. However, currently, the best performance is achieved by deep neural networks that can learn complex detection mechanisms from the data. Originally, convolutional neural network architectures were proposed but recently also transformer-based architectures have gained popularity in the pedestrian detection task, DETR [6] being a popular example. Convolutional architectures are typically further divided into one-stage and two-stage detectors. One-stage detectors predict the bounding box coordinates and related class probabilities in a single forward pass. Common example is the YOLO family [7]. Two-stage detectors first propose regions of interest and then classify and refine the proposals. Common examples are RCNN [8] and its predecessors Fast RCNN [9] and Faster RCNN [10]. To support the development of the pedestrian detection models variety of datasets have been presented over the years, e.g. Caltech Pedestrian Dataset [11], CityPersons dataset [12], KITTI [13] and WiderPerson [14].
II-B Detection in low-light conditions
@@
In low-light conditions, pedestrian detection from conventional visible-spectrum RGB images is significantly more challenging as the lack of visible light leads to low contrast, loss of color information, and severe noise effects. To improve low-light RGB detection, image enhancement methods which aim to restore the low-light image before feeding it to a standard detector, have been proposed [15, 16]. Improved performance can be achieved by optimizing the enhancement for the object detection task specifically by end-to-end training [17, 18]. However, superior performance can be achieved by leveraging additional sensors like a thermal camera or lidar that provide measurements which are much less affected by the lighting conditions. Lidar-camera fusion for object detection has been extensively researched and [19, 20] are notable examples including low-light conditions in the experiments. RGB-thermal fusion has also been explored in multiple studies. A popular baseline solution is provided in [21] and recent work presenting a more advanced fusion strategy in [22]. In [23, 24] teacher-student architecture is proposed where the teacher has access to RGB and thermal images while the students can only rely on RGB images to train a reliable RGB detector for low-light conditions. While most of the available data is from daytime, some datasets addressing the low-light conditions have been published. KAIST [21], LLVIP [25], FLIR-ADAS [26] and CVC-14 [27] provide RGB and thermal images, and Nightowls [28] only provides RGB images.
II-C Automated learning
@@
Detection models perform poorly outside their training domain, and manual annotation is one of the main bottlenecks for training data generation. Thus, research efforts have also focused on ways to reduce or remove the need for manual annotations. A common approach is unsupervised domain adaptation, where the model is adapted from a source domain to a target domain and annotations are only available in the source domain. In [29] RGB-thermal pedestrian detection model was adapted from the KAIST dataset domain to the CVC-14 dataset domain in unsupervised manner with performance close to supervised methods. Also in [30] unsupervised domain adaptation yields comparable results to supervised methods when moving between the CVC-14, KAIST and FLIR ADAS datasets. Another way to reduce the need for labeling is to transfer the labels between sensors. If reliable detections are available in one sensor frame they can be moved to other sensor frames. In [31] bounding boxes from pre-trained RGB detector were transfered to the corresponding thermal frames to train thermal detection models without the need for manual labeling. Arguably for low-light pedestrian detection transfering the labels in opposite direction would be lucrative. Reliable detections on thermal images could be transferred to the corresponding RGB frames to train a RGB prediction model.
@@ III Methods
This paper presents an automatic pedestrian labelling method for RGB image data, leveraging corresponding IR images (Fig.1). The implementation is aimed at enhancing pedestrian detection in low-light conditions. The proposed architecture is presented in Section III-A. The KAIST dataset used in the paper is described in Section III-B, and the implementation details are presented in Section III-C.
@@ III-A2 Label transfer
In the KAIST dataset [21] used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
@@ III-B Data
The proposed autolabelling method is evaluated with the KAIST dataset. KAIST includes $\sim$95k samples in total, where each sample includes an aligned IR-RGB frame pair and manual ground truth annotation for the cyclist, person, and people classes. The people class is used if individual persons cannot be distinguished from a group of people. In our experiments, we excluded the samples that include cyclist or people labels, yielding $\sim$91.8k samples. This exclusion was done due to the three classes being too closely related to each other, resulting in extremely poor performance for any models trained on the RGB low-light portion of the dataset. Of the remaining images, $\sim$52.7k samples are from daytime (sequences 00, 01, 02, 06, 07, 08) and $\sim$23.7k samples from lowlight conditions (sequences 03, 04, 05, 09, 10, 11). The IR models were trained with all of the daytime sequences except sequence 02, which was retained for testing, resulting in $\sim$51k training images and $\sim$7.3k testing images. The low-light RGB models were trained with all of the low-light sequences except sequence 09, which was retained for testing, resulting in $\sim$26.7k training images and $\sim$3.5k testing images.
@@ III-C Evaluation
We chose to evaluate our pipeline with the YOLOv11[32], RCNN [10], and DETR[33] architectures. The specific versions chosen were YOLOv11l, Faster RCNN, and RF-DETR. All three models were trained on the daytime IR frames for 20 epochs with the default hyperparameters. These fine-tuned versions of the models are referred to as IR-YOLO, IR-RCNN, and IR-DETR. For all three models, the models were initialized with COCO [34] pretrained weights. The fine-tuned IR models were used for autolabelling the low-light portion of the dataset. To evaluate how the label confidence threshold affects the final results, two label sets were generated with each model, using 0.25 and 0.5 confidence thresholds. Additionally, the models were also trained with ground truth labels, allowing for performance comparisons between models trained on generated and ground truth labels.
@@
Same YOLO, DETR, and RCNN variants were also used for the low-light RGB pedestrian detection. Each of the three models was trained on its IR version labels generated with 0.25 and 0.5 confidence thresholds, and the ground truth labels for comparison. This resulted in a total of 9 low-light RGB pedestrian detection models. Similar to the IR models, each model was trained for 20 epochs with the default hyperparameters using COCO pretrained weights as the starting point.
@@ IV Results
The number of labels generated with each of the IR detection models at 0.25 and 0.5 confidence thresholds are presented in Table I. The KAIST ground truth label count is included for comparison. IR-DETR with 0.25 confidence generated the highest number of labels, and IR-YOLO with 0.5 confidence generated the least amount of labels. To assess the accuracy of the generated labels, standard object detection metrics Precision(P), Recall(R), F1, mAP@50,mAP@50-95 and Log Average Miss Rate (LAMR) are also presented at both confidence levels (Table II). IR-DETR had the highest overall performance on the mAP and LAMR metrics, and the labels generated with IR-DETR at 0.5 confidence had the highest F1 score. YOLO had the second-highest and RCNN the lowest performance on the mAP and LAMR metrics. For RCNN and YOLO 0.25 confidence threshold yielded a higher F1 score than a 0.5 confidence threshold. The precision-recall curves for the IR models are presented in Figure 2.
@@
The best performance overall was achieved by DETR trained on generated labels with a confidence value of 0.25. The precision-recall curves for each RGB model variant are presented in Figure 3. Each architecture exhibits quite similar precision-recall tradeoffs regardless of the training labels used, while the differences between architectures are significantly larger.
@@
0.25 Conf 0.50 Conf GT labels GT IR GT RGB
Prediction model
DETR
@@
YOLO
RCNN
@@ V Discussion
In this paper, an automated labelling method for low-light pedestrian detection leveraging corresponding IR and RGB images was presented. YOLOv11l, RF-DETR, and Faster RCNN architectures were tested for both the IR image-based autolabelling and lowlight RGB detection, trained with the autolabeled data. Training with the autolabels yielded higher mAP@50 and LAMR than training with the ground truth labels in 5 out of 6 cases and higher mAP@50-95 in all cases.
@@
Training with the autolabels yielded better results than training with the ground truth labels as shown in Table III and demonstrated on the testing data in Figure 4. This difference in performance could be attributed to the hard cases in ground truth labels, where pedestrians are extremely poorly visible in the RGB images due to occlusion and poor illumination, with examples of such cases shown in Figure 5. Such samples can cause the model to overfit to the training data, as generalizable detection features are not found, resulting in lower overall performance. Another possible explanation is the difficulty of the low-light RGB detection task. The mAP scores of the IR models are up to two times higher compared to the RGB models. Likely, the accuracy of the autolabels is not the main bottleneck, but instead the limitations of RGB detection in low-light conditions.
In our experiments, the model architecture was found to play a major role in the final performance. All IR models were trained on the same data, DETR outperforming YOLO and RCNN with a significant margin (Table II). A similar pattern was observed when low-light RGB models were trained on the ground truth, DETR being the best-performing model (Table III). These results underline the importance of model selection, as DETR was able to adapt to both infrared and low-light conditions significantly better using the same data as the other models.
Another parameter explored in this paper was the autolabelling detection confidence threshold. When the confidence threshold was lowered from 0.5 to 0.25, Recall increases and Precision decreases for all models, while the F1 score increases for IR-YOLO and IR-RCNN and maintains a similar value for IR-DETR. (Table II). The low-light RGB models trained on the autolabels generated with a confidence threshold of 0.25 performed the best in all cases (Table III). There seemed to be a correlation between the F1 score of the autolabels and the performance of the resulting RGB prediction model.
Overall, all of the low-light RGB models have a low recall rate, with the highest being 0.369 for DETR trained on the IR-DETR labels with 0.25 confidence. This highlights both the need for more quality training data and the potential upside from sensor fusion approaches. Camera-based detection is highly affected by lighting conditions, while less lighting-sensitive sensors can provide reliable additional sources of information. Finally, this approach does not pose computation time limitations, allowing for the usage of ensemble models for autolabelling.

### v3

@@ I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian-related traffic accidents [ped_Det]. According to the World Health Organization (WHO) [who2023traffic], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors, such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions, nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [moore2024nightdriving] on the data provided by the Transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
Figure 1: Autolabelling pipeline, using unlabelled low-light IR-RGB image pairs as input, the system outputs labelled RGB images, enabling the training of low-light pedestrian detection models
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in low-light conditions, as RGB model accuracy is highly affected by lighting [obj_det_issues]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared (IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal, we leverage the lighting invariance of IR cameras as visible light has minimal effect on IR radiation. In our approach, we utilise the benefits of IR images to detect pedestrians in IR and then transfer the labels to the corresponding RGB image. Our proposed approach relies on aligned IR and RGB images, which can be conveniently collected with a research vehicle equipped with a suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work. The proposed pipeline has not previously been implemented for autolabelling tasks and shows a range of positive initial results. The autolabels generated using the proposed pipeline achieved the goal of performing on par or outperforming the ground truth labels in many cases, in key metrics such as mAP@50, mAP@50-95, and LAMR. Consequently, the pipeline provides all of the upside gained from autolabelling by eliminating human effort completely while resulting in detection models on par with those trained on manually labelled data.
@@ II-A Pedestrian Detection
Early pedestrian detection approaches utilise hand-crafted features such as Histogram of Oriented Gradients (HOG) [10YEARSOF-peddet]. However, currently, the best performance is achieved by deep neural networks that can learn complex detection mechanisms from the data. Originally, convolutional neural network architectures were proposed but recently, transformer-based architectures have also gained popularity in the pedestrian detection task, DETR [carion2020end] being a popular example. Convolutional architectures are typically further divided into one-stage and two-stage detectors. One-stage detectors predict the bounding box coordinates and related class probabilities in a single forward pass. A common example is the YOLO model family [redmon2016you]. Two-stage detectors first propose regions of interest and then classify and refine the proposals. Common examples are RCNN [girshick2014rich] and its predecessors Fast RCNN [girshick2015fast] and Faster RCNN [rcnn_citation]. To support the development of the pedestrian detection models variety of datasets have been presented over the years, e.g. Caltech Pedestrian Dataset [dollar2009pedestrian], CityPersons dataset [CityPersons-DATASET], KITTI [KITTI-DATASET], and WiderPerson [widerperson_dataset].
II-B Pedestrian detection in low-light conditions
@@
In low-light conditions, pedestrian detection from conventional visible-spectrum RGB images is significantly more challenging as the lack of visible light leads to low contrast, changes in feature visibility, and severe noise effects. To improve low-light RGB detection, image enhancement methods which aim to restore the low-light image before feeding it to a standard detector have been proposed [cai2023retinexformer, wei2018deep]. Improved performance can be achieved by optimising the enhancement for the object detection task, specifically by end-to-end training [liu2022image, guo2021dynamic]. However, superior performance can be achieved by leveraging additional sensors like an IR camera or lidar that provide measurements less affected by the lighting conditions. Lidar-camera fusion for object detection has been extensively researched, and [zhang2022robust, bai2022transfusion] are notable examples, including low-light conditions in the experiments. IR-RGB fusion has also been explored in multiple studies. A popular baseline solution is provided in [KAIST_dataset] and recent work presenting a more advanced fusion strategy in [zhang2024tfdet]. In [liu2021deep, kruthiventi2017low] teacher-student architecture is proposed where the teacher has access to RGB and IR images, while the students can only rely on RGB images to train a reliable RGB detector for low-light conditions. While most of the available data is from daytime, some datasets addressing the low-light conditions have been published. KAIST [KAIST_dataset], LLVIP [LLVIP_dataset], FLIR-ADAS [FLIRADAS_dataset] and CVC-14 [cvc14_dataset] provide RGB and IR images, and NightOwls [nightowls_dataset] only provides RGB images.
II-C Automated labelling
@@
Detection models perform poorly outside their training domain, and manual annotation is one of the main bottlenecks for training data generation. Thus, research efforts have also focused on ways to reduce or remove the need for manual annotations. A common approach is unsupervised domain adaptation, where the model is adapted from a source domain to a target domain, and annotations are only available in the source domain. In [domain_adaptation_autolabelling] IR-RGB pedestrian detection model was adapted from the KAIST dataset domain to the CVC-14 dataset domain in an unsupervised manner with performance close to supervised methods. Also in [usupervised_ped_det_ir-thermal_autolabelling], unsupervised domain adaptation yields comparable results to supervised methods when moving between the CVC-14, KAIST and FLIR ADAS datasets. Another way to reduce the need for labelling is to transfer the labels between sensors. If reliable detections are available in one sensor frame, they can be moved to other sensor frames. In [canton2024automatic], bounding boxes from a pre-trained RGB detector were transferred to the corresponding IR frames to train IR detection models without the need for manual labelling. The potential of the opposite label transfer method, from IR to RGB, was explored during the creation of the LLVIP [LLVIP_dataset] dataset. The dataset consists of aligned IR-RGB frames, and due to pedestrians being more distinct in IR, the manual labelling was performed in IR, with the subsequent step of transferring the labels to RGB. The groundwork done in LLVIP indicates the potential in labelling IR frames and using those labels for RGB detection tasks. The goal of this research is to explore the idea of fully automating the IR to RGB labelling process by autolabelling data in IR and then transferring the labels to corresponding RGB images. This enables scalable data collection and extensive training of low-light pedestrian detection models for RGB images.
@@ III Methods
This paper presents an automatic pedestrian labelling method for RGB image data, leveraging corresponding IR images (Fig.1). The implementation is aimed at enhancing pedestrian detection in low-light conditions. The proposed architecture is presented in Section III-A. The KAIST Multispectral Pedestrian Detection Benchmark dataset[KAIST_dataset] used in the paper is described in Section III-B, and the implementation details are presented in Section III-C.
@@ III-A2 Label transfer
In the KAIST dataset [KAIST_dataset] used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
@@ III-B Data
The proposed autolabelling method is evaluated with the KAIST dataset. KAIST includes $\sim$95k samples in total, where each sample includes an aligned IR-RGB frame pair and manual ground truth annotation for the cyclist, person, and people classes. The people class is used if individuals cannot be distinguished from a group of people. In our experiments, we excluded the samples that include cyclist or people labels, yielding $\sim$91.8k samples. This exclusion was done due to the three classes being too closely related to each other, resulting in extremely poor performance for any models trained on the RGB low-light portion of the dataset due to classification errors. Of the remaining images, $\sim$52.7k samples are from daytime (sequences 00, 01, 02, 06, 07, 08) and $\sim$23.7k samples from lowlight conditions (sequences 03, 04, 05, 09, 10, 11). The IR models were trained with all of the daytime sequences except sequence 02, which was retained for testing, resulting in $\sim$51k training images and $\sim$7.3k testing images. The low-light RGB models were trained with all of the low-light sequences except sequence 09, which was retained for testing, resulting in $\sim$26.7k training images and $\sim$3.5k testing images.
@@ III-C Evaluation
We chose to evaluate our pipeline with the YOLOv11 [yolov11_ultralytics_paper], RCNN [rcnn_citation], and DETR [RF-DETR_2025_citation] architectures. The specific versions chosen were YOLOv11l, Faster RCNN, and RF-DETR. All three models were trained on the daytime IR frames for 20 epochs with the default hyperparameters. These fine-tuned versions of the models are referred to as IR-YOLO, IR-RCNN, and IR-DETR. For all three models, the models were initialised with COCO [COCO_dataset] pretrained weights. The fine-tuned IR models were used for autolabelling the low-light portion of the dataset. To evaluate how the label confidence threshold affects the final results, two label sets were generated with each model, using 0.25 and 0.5 confidence thresholds. Additionally, the models were also trained with ground truth labels, allowing for performance comparisons between models trained on generated and ground truth labels.
@@
Same YOLO, DETR, and RCNN variants were also used for the low-light RGB pedestrian detection for reproducibility purposes. Each of the three models was trained on its IR version labels generated with 0.25 and 0.5 confidence thresholds, and the ground truth labels for comparison. This resulted in a total of 9 low-light RGB pedestrian detection models. Similar to the IR models, each model was trained for 20 epochs. Training was stopped at 20 epochs as validation metrics plateaued, and additional epochs did not result in measurable improvement. For the training, the default hyperparameters were used, with COCO-pretrained weights serving as the starting point.
@@ IV Results
The number of labels generated with each of the IR detection models at 0.25 and 0.5 confidence thresholds is presented in Table I. The KAIST ground truth label count is included for comparison. IR-DETR with 0.25 confidence generated the highest number of labels, and IR-YOLO with 0.5 confidence generated the fewest number of labels. To assess the accuracy of the generated labels, standard object detection metrics Precision(P), Recall(R), F1, mAP@50, mAP@50-95 and Log Average Miss Rate (LAMR) are also presented at both confidence levels (Table II). IR-DETR had the highest overall performance on the mAP and LAMR metrics, and the labels generated with IR-DETR at 0.5 confidence had the highest F1 score. YOLO had the second-highest and RCNN the lowest performance on the mAP and LAMR metrics. For RCNN and YOLO 0.25 confidence threshold yielded a higher F1 score than a 0.5 confidence threshold. The precision-recall curves for the IR models are presented in Figure 2.
@@
The best performance overall was achieved by DETR trained on generated labels with a confidence value of 0.25. The precision-recall curves for each RGB model variant are presented in Figure 3. Each architecture exhibits quite similar precision-recall tradeoffs regardless of the training labels used, while the differences between architectures are significantly larger. These results showcase the benefit of autolabelling as they performed on par or better than manually generated labels while requiring no human effort or input.
@@
IR-DETR 0.25 Conf IR-DETR 0.50 Conf GT labels GT IR GT RGB
Prediction model DETR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
IR-YOLO 0.25 Conf IR-YOLO 0.50 Conf GT labels GT IR GT RGB
YOLO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
IR-RCNN 0.25 Conf IR-RCNN 0.50 Conf GT labels GT IR GT RGB
RCNN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ V Discussion
This paper presented an automated labelling pipeline for low-light pedestrian detection using IR-RGB image pairs. For IR autolabelling and RGB pedestrian detection, YOLOv11l, RF-DETR, and Faster RCNN baseline models were used. The evaluation showed that training on the generated autolabels yields higher mAP@50 and lower LAMR compared to ground truth labels provided in KAIST across all tests and in 5 out of 6 for mAP@50-95, achieving the goal of matching human-generated ground truth in performance without the need for human input.
@@
Training with the autolabels yielded better results than training with the ground truth labels, as shown in Table III and demonstrated on the testing data in Figure 4, consequently exceeding the initial goal of matching ground truth performance. This difference in performance could be attributed to two factors: the hard cases in ground truth labels and the general difficulty of low-light pedestrian detection. In the ground truth labels, examples where pedestrians are extremely poorly visible or heavily occluded exist, with samples shown in Figure 5. These samples can potentially act as label noise that are technically correct but lack features and visual cues, consequently lowering model performance. As object detection in low-light conditions often results in lowered accuracy, the combination of these two factors could be the reason for the performance drop. The same effect is not observed in the autolabels, as heavily occluded and partially visible pedestrians are ignored, due to the autolabelling model not detecting them as pedestrian instances, consequently filtering out extremely hard cases.
In our experiments, another key finding was the effect of model architecture on the final output. In all cases where the three baseline models were trained on the same labels, we observe that DETR performs best, followed by YOLO and lastly by RCNN. This is observed in IR model training (Table II) and in the case where ground truth labels are used (Table III) during RGB model training. Furthermore, the effect of the model is highlighted as all DETR models performed better than any other model-label combination on the low-light RGB detection task. This observation shows the importance of appropriate model selection.
The results presented in Table III not only show the importance of quality labels and appropriate model selection, but also the challenge of low-light object detection as a whole. Of the presented models, regardless of model-label pair, all cases did not produce a model that can be reliably used as a standalone detection tool. This is concluded due to the best recall being 0.369, indicating that the models are not capable of reliably finding all pedestrians in the images. This highlights the difficulty of the RGB pedestrian detection task in low-light conditions. However, as RGB cameras are readily available in vehicles, the detection capabilities are still valuable for different intelligent systems. Larger datasets enable the training and development of more accurate models for the detection task. Our work directly contributes to this, as the proposed autolabelling method can be used for collecting extensive datasets.
Another parameter explored in this paper was the autolabelling detection confidence threshold. In all cases when lowering confidence from 0.5 to 0.25, Precision decreased, and Recall increased, F1 increased for IR-YOLO and IR-RCNN and decreased for IR-DETR (Table II). Additionally, when autolabels were generated in all cases, the models running on 0.25 confidence produced significantly more labels compared to 0.5 (Table I). When low-light pedestrian detection models were trained (Table III), all models performed better on mAP@50, mAP@50-95, and LAMR. This shows that additional labels generated during 0.25 confidence automatic labelling had a positive effect on subsequent training, resulting in another parameter to be further explored and verifying the need for more overall data.